# Fully Adaptive Zeroth-Order Method for Minimizing Functions with Compressible Gradients

### v2

@@ 3.2 Oracle Complexity Assuming Compressible Gradients
Notice that $J=\left\lceil\log_{2}\left(\frac{n}{bs_{0}\ln(n)}\right)\right\rceil$ is the smallest non-negative integer such that $m_{J}=\lceil bs_{J}\ln(n)\rceil\geq n$. Therefore,
@@
are the only possible sensing matrices used in Algorithm 1. In this subsection we will consider the following additional assumptions:
@@
Lemma 3.7.
@@
Then, it follows from (38), (40) and Corollary 2.6 that (14) holds for $x=x_{k}$ and $g=g_{k}^{(j^{*})}$. Thus, by (39) and Lemma 2.7 we would have
@@
Lemma 3.7 implies the following upper bounds on $s_{k}$ and $\sigma_{k}$.
Lemma 3.8.
@@
Thus, combining (42) and the second inequality in Lemma 3.8 we see that (41) is true. ∎
Corollary 3.10.
@@
$FE_{T(\epsilon)}$$\leq$ $4\left[1+\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right]\times$ (43)
@@
Proof.
Let us consider the $k$th iteration of Algorithm 1. By Lemma 3.7 and Assumption A4, we have $m_{j}=bs_{j}\ln(n)<n$ for $j=0,\ldots,j_{k}$. Then, for $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $(m_{j}+1)$ function evaluations. Moreover, checking the acceptance condition (24) requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by
$\sum_{j=0}^{j_{k}}(m_{j}+2)$$\leq$ $2\sum_{j=0}^{j_{k}}bs_{j}\ln(n)\leq 2(1+j_{k})s_{j_{k}}\ln(n)$
$\leq$ $2\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$
@@
$FE_{T(\epsilon)}$$\leq$ $2\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$ (44)
@@
When $s_{0}\leq s(\theta,p)$, it follows from Corollary 3.10 that
@@
$FE_{T(\epsilon)}\leq\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right).$
This upper complexity bound relies on the assumption that $Z^{(j^{*})}$ satisfy the $4s_{j^{*}}$-RIP with $\delta_{4s_{j^{*}}}(Z^{(j^{*})})<0.22665$ (Assumption A5). In view of Lemma 2.3 and our construction of matrices $Z^{(j)}$, if A1-A4 hold and constant $b$ is sufficiently large, A5 holds with with high probability. Consequently, with high probability, Algorithm 1 needs no more than $\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\,\cdot\,)$.
@@ 4 Illustrative Numerical Experiments
In the following experiments, we compare ZORO-FA to a representative sample of deterministic DFO algorithms, namely Nelder-Mead [20], DFQRM [11], and ZORO [6]. We avoid methods which incorporate curvature, or “second-order” information such as CARS [12], NEWUOA [21, 27] or the derivative-free variation of L-BFGS studied in [24], noting that such functionality could be added to ZORO-FA fairly easily.
@@ 4.1 Sparse Gradient Functions
We test ZORO-FA on a challenging test function known to exhibit gradient sparsity, namely the max-s-squared function ($f_{ms}$) [5, 23] defined as
@@
$\pi(\cdot)$ is the permutation ordering the coordinates of $x$ from largest to smallest in magnitude. This function is challenging as, while $\nabla f_{ms}$ is $p$-compressible (in fact, sparse), the location of the largest-in-magnitude entries $\nabla f_{ms}$ changes frequently. Moreover, $\nabla f_{ms}$ is not Lipschitz continuous. We take $n$, the dimension of the function, to be $1000$ and $s=30$.
@@
We tuned the parameters of each algorithm to the best of our ability; see Table 1 for exact values used. In all trials $x_{0}$ is a random draw from the multivariate normal distribution with mean 00 and covariance $10I$. Each algorithm is given a query budget of $500(n+1)$.
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta\ddagger$ $\sigma_{0}$ $\sigma_{\min}$ min. simplex size
@@
ZORO $1$ $0.5$ $30$
adaZORO $1$ $0.5$ $30$
ZORO-FA $1$ $20$ $0.01$ $0.25$ $1$
Nelder-Mead - - - - - - - $0.001$
Table 1: Parameters used in sparse benchmarking experiments of Section 4.1. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, and the initial target sparsity of adaZORO and ZORO-FA (which both dynamically adjust this parameter). $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning.
@@ Results.
Figure 1 displays a typical trajectory, as well as box plots for the run time over ten trials. Note that the only source of randomness in each trial is the initial point $x_{0}$. Unsurprisingly, ZORO and adaZORO perform well, as they are given the exact sparsity as an input and have hand-tuned step-sizes. While Nelder-Mead and DFQRM struggle, ZORO-FA makes steady progress.
Figure 1: Left: Typical trajectories for all five algorithms. Note the $x$-axis displays the cumulative number of queries divided by $n+1$. Right: Wall-clock run time.
@@ 4.2 Moré-Garbow-Hillstrom Test Functions
Next, we benchmark ZORO-FA on a subset of the well-known Moré-Garbow-Hillstrom (MGH) suite of test problems [16].
@@
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta$ $\sigma_{0}$ $\sigma_{\min}$ min. simplex size
DFQRM - - - $10^{-5}$ 00$\ddagger$0 ‡ $1$ $0.01$
ZORO $1$ $0.005$ $0.1n$
adaZORO $1$ $0.005$ $0.1n$
ZORO-FA $1$ $0.1n$ $0.01$ $0.25$ $1$
Nelder-Mead - - - - - - - $0.001$
Table 2: Parameters used in the Moré-Garbow-Hillstrom benchmarking experiments of Section 4.2. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, and the initial target sparsity of adaZORO and ZORO-FA (which both dynamically adjust this parameter). $n$ refers to the problem size. $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning.
@@
We present the results as data profiles [17] for various tolerances $\tau$. Let $\mathcal{P}$ denote the test problem set. A point $(\alpha,y)$ on a curve in a data profile indicates that a fraction of $y|\mathcal{P}|$ problems were solved, to tolerance $\tau$, given $(n+1)\times\alpha$ function evaluations. We say an algorithm solves a problem to tolerance $\tau$ if it finds an iterate $x_{k}$ satisfying
@@ 4.3 Discussion
ZORO-FA outperforms DFQRM and Nelder-Mead on functions exhibiting gradient sparsity, and is in turn outperformed by ZORO and adaZORO. The latter statement is not surprising, as ZORO and adaZORO are given the exact gradient sparsity parameter and a hand-tuned step-size. The situation is reversed for the MGH test functions; ZORO and adaZORO struggle — presumably as they are unable to adjust their step sizes — while Nelder-Mead, DFQRM, and ZORO-FA perform well. It was surprising, at least to us, that Nelder-Mead performed so well. We note however that this performance came at high computational expense, with Nelder-Mead typically taking 10–100 times longer to run (see the right panel in Figure 1).
@@ 4.4 Hidden Gradient Compressibility
In contrast to Section 4.1, the MGH functions of Section 4.2 are not explicitly designed to possess sparse or compressible gradients. Nonetheless we observe that ZORO-FA works well, and is frequently able to make progress with $s_{k}\ll n$. To investigate this further, we plot the magnitudes of the entries of $\nabla f(x)$, sorted from largest to smallest, for several MGH test functions at $20$ randomly selected $x$. Letting $|\nabla f(x)|_{(j)}$ denote the $j$-th largest-in-magnitude component of $\nabla f(x)$, we plot the mean of $|\nabla f(x)|_{(j)}$ over all $20$ values of $x$, as well as the minimum and maximum values of $|\nabla f(x)|_{(j)}$ observed. The results are displayed in Figure 3. As is clear, some MGH test functions (e.g., rosex) exhibit gradient compressibility, while others (e.g., trig) do not. This further emphasizes the practical need for optimization methods which can take advantage of the presence of gradient sparsity, yet are robust to its absence.
Figure 3: Gradient magnitude profiles for two MGH functions: rosex (Left), and trig (Right). The solid line represents the mean (over 20 trials), while the shading represents the minimum and maximum values of gradient component magnitudes observed.

### v3

@@ 3.2 Oracle Complexity Assuming Compressible Gradients
In this subsection, we analyze Algorithm 1 under the additional assumption that the gradients of $f(\,\cdot\,)$ are compressible (Assumption A3). Note that
$J\equiv\left\lceil\log_{2}\left(\frac{n}{bs_{0}\ln(n)}\right)\right\rceil$
is the smallest positive integer such that $m_{J}=\lceil bs_{J}\ln(n)\rceil\geq n$. Therefore,
@@
are the only possible sensing matrices used in Algorithm 1. Moreover, since $s_{0}$ is chosen such that $\left\lceil bs_{0}\ln(n)\right\rceil\leq n/4$, we have $J\geq 2$. In this subsection we will consider the following additional assumptions:
@@
Remark 4
Assumption 4 implicitly requires $s(\theta,p)$ to be sufficiently small. In view of (12), this means that the gradients should be $p$-compressible (Assumption A3), with $p$ sufficiently close to zero. Moreover, by Lemma 3, Assumption A5 holds with high probability when $b$ in Step 0 is chosen sufficiently large. This fact is clarified in Corollary 4.
Lemma 10
@@
Then, it follows from (39), (41) and Corollary 1 that (15) holds for $x=x_{k}$ and $g=g_{k}^{(j^{*})}$. Thus, by (40) and Lemma 6 we would have
@@
Lemma 10 implies the following upper bounds on $s_{k}$ and $\sigma_{k}$.
Lemma 11
@@
Thus, combining (43) and the second inequality in Lemma 11 we see that (42) is true.∎
Corollary 3
@@
$FE_{T(\epsilon)}$$\leq$ $4b\left[1+\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right]\times$ (44)
@@
Proof
Let us consider the $k$th iteration of Algorithm 1. By Lemma 10 and Assumption A4, we have $m_{j}=bs_{j}\ln(n)<n$ for $j=0,\ldots,j_{k}$. Then, for $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $(m_{j}+1)$ function evaluations. Moreover, checking the acceptance condition (25) requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by
$\sum_{j=0}^{j_{k}}(m_{j}+2)$$\leq$ $2\sum_{j=0}^{j_{k}}bs_{j}\ln(n)\leq 2b(1+j_{k})s_{j_{k}}\ln(n)$
$\leq$ $2b\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$
@@
$FE_{T(\epsilon)}$$\leq$ $2b\left[1+\left\lceil\max\left\{1,\log_{2}\left(\dfrac{s(\theta,p)}{s_{0}}\right),\log_{2}\left(\dfrac{(\theta+1)^{2}}{(1-2\theta)}\dfrac{L}{\sigma_{0}}\right)\right\}\right\rceil\right]\times$ (45)
@@
When $s_{0}\leq s(\theta,p)$, it follows from Corollary 3 that
@@
$FE_{T(\epsilon)}\leq\mathcal{O}\left(s(\theta,p)\ln(n)\epsilon^{-2}\right).$ (46)
Corollary 4
Suppose Assumptions A1–A4 hold, and consider the functions $c_{1}(\cdot)$ and $\gamma(\cdot)$ defined in (2) and (3), respectively. If $b\geq c_{1}(0.22664)$, then, with probability at least
$1-2e^{-\gamma(0.22664)m_{j^{*}}},$ (47)
Algorithm 1 requires no more than $\mathcal{O}\bigl{(}s(\theta,p)\ln(n)\epsilon^{-2}\bigr{)}$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\cdot)$.
Proof
If Assumption A5 also holds, then the stated complexity bound follows from Corollary 3. Since $b\geq c_{1}(0.22664)$, it follows from Lemma 3 that, with probability at least (47), the sensing matrix $Z^{(j^{*})}$ satisfies the $4s_{j^{*}}$-RIP with $\delta_{4s_{j^{*}}}(Z^{(j^{*})})<0.22665$. This implies that Assumption A5 holds with probability at least (47). Therefore, the stated complexity bound holds with the same probability.
@@ 4 Illustrative Numerical Experiments
In the following experiments, we compare ZORO-FA to a representative sample of deterministic DFO algorithms, namely Nelder-Mead NM , DFQRM grapiglia2024quadratic , the zeroth-order variant kozak2023 of stochastic subspace descent Kozak (henceforth: SSD), direct search in reduces spaces Roberts (henceforth: DS-RS), and ZORO and adaZORO cai2022zeroth . We avoid methods which incorporate curvature, or “second-order” information such as CARS kim2021curvature or the derivative-free variation of L-BFGS studied in shi2023numerical , noting that such functionality could be added to ZORO-FA fairly easily. We also do not compare against model-based approaches which use a quadratic model, such as DFBGN Cartis or NEWUOA powell2006newuoa ; Zhang_2023 , as these offer different trade-offs between run-time, query complexity, and solution quality.
@@ 4.1 Sparse Gradient Functions
We test ZORO-FA on two challenging test functions known to exhibit gradient sparsity, namely the max-s-squared function ($f_{ms}$) cai2022one ; slavin2022adapting and a variant of Nesterov’s ‘worst function in the world’ Kozak ; nesterov2013 ($f_{N,\lambda,s}$). These functions are defined as
@@
$f_{N,\lambda,s}(x)$ $=\frac{\lambda}{8}\left(x_{1}^{2}+\sum_{i=1}^{s}\left(x_{i}-x_{i+1}\right)^{2}+x_{s}^{2}\right)-\frac{\lambda}{4}x_{1}.$
The function $f_{ms}$ is challenging as, while $\nabla f_{ms}$ is $p$-compressible (in fact, sparse), the location of the largest-in-magnitude entries $\nabla f_{ms}$ changes frequently. Moreover, $\nabla f_{ms}$ is not Lipschitz continuous. The function $f_{N,\lambda,s}$ is a variant of a well-known pathological example used to prove tightness of convergence rate bounds in convex optimization. Note that $\nabla f_{N,\lambda,s}$ is Lipschitz continuous with Lipschitz constant $\lambda/2$. For both functions, we take $n$, the dimension of the function, to be $1000$ and $s=30$. For $f_{N,\lambda,s}$ we take $\lambda=8$ as in Kozak .
@@
We used the default parameters given in the code for DFQRM, as these worked well. For DS we used the parameters given in Section 4 of Roberts for the Gaussian polling direction, $r=1$, setting. [^3] That is, at each iteration, we query $f$ at $f(x_{k}+z)$ and $f(x_{k}-z)$ where $z$ is a Gaussian random vector as this setting appears to work best. That is, with notation as in Roberts , we take $\alpha_{0}=1.0,\alpha_{\max}=1000,\gamma_{\mathrm{inc}}=2,\gamma_{\mathrm{dec}}=0.5$, and we terminate whenever $\alpha_{k}<10^{-6}$. For SSD there did not appear to be a ‘best’ choice of subspace dimension in the results of Kozak . So, we take this to be equal to $s$ in our experiments. [^4] We experimented with other values, and did not see significantly different performance.. Three algorithms require a step-size parameter: ZORO, adaZORO, and SSD, which should theoretically be equal to or proportional to the inverse of the Lipschitz constant of the objective function. In order to guarantee convergence for both $f_{ms}$ and $f_{N,\lambda,s}$, we set the step size to the more pessimistic value [^5] Although $f_{ms}$ is not everwhere differentiable, it has a Lipshitz constant of $2$ wherever it is differentiable. of $1/\lambda$. Finally, both ZORO and adaZORO are given the exact sparsity. To make the setting more challenging for ZORO-FA, we deliberately provide it with an underestimate of the true sparsity as $s_{0}$, namely $20$. See Table 1 for the values of the major parameters used. In all trials $x_{0}$ is a random draw from the multivariate normal distribution with mean 00 and covariance $10I$. Each algorithm is given a query budget of $350(n+1)$.
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta\ddagger$ $\sigma_{0}\lozenge$ $\sigma_{\min}^{\star}$ samp. radius
@@
ZORO $1$ $1/8$ $30$ $10^{-4}$
adaZORO $0.5$ $1/8$ $30$ $10^{-4}$
ZORO-FA $1$ $20$ $10^{-5}$ $0.25$ $2.5$
Nelder-Mead - - - - - - $0.001$
SSD - $\frac{s_{0}}{8n}$ $30$ $10^{-4}$
DS-RS - - $1$ 1 $10^{-6}$
Table 1: Parameters used in sparse benchmarking experiments of Section 4.1. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, the initial target sparsity of adaZORO and ZORO-FA, and the subspace dimension of SSD and DS. $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning. $\lozenge$: For DS, this is the initial step size ($\alpha_{0}$ in Roberts ). $\star$: In Nelder-Mead this is the minimum simplex size; in DS-RS this is threshold for $\alpha_{k}$ at which we terminate.
@@ Results.
Figure 1 displays typical trajectories for both functions. We present just a single trajectory as we observed very little variation between runs. As is clear, ZORO-FA substantially outperforms all other algorithms. Although ZORO and adaZORO descend quickly at the beginning—likely due to the fact they are given the exact sparsity as input—their progress quickly stalls out as they are unable to decrease their step size or sampling radius. Finally, we note that all algorithms save Nelder-Mead had similar run times—around 1010101050505050 seconds. Nelder-Mead typically took two orders of magnitude more time to run.
Figure 1: Left: Typical trajectories on $f_{ms}$. adaZORO fails when it is unable to find a gradient approximation satisfying its acceptance criterion. DS-RS fails when $\alpha_{k}$ fall below $10^{-6}$. Right: Typical trajectories on $f_{N,\lambda,s}$. Note the $x$-axis displays the cumulative number of queries divided by $n+1$.
@@ 4.2 Moré-Garbow-Hillstrom Test Functions
Next, we benchmark ZORO-FA on a subset of the well-known Moré-Garbow-Hillstrom (MGH) suite of test problems more1981testing . For simplicity, we do not include SSD or DS-RS in this experiment.
@@
We present the results as data profiles more2009benchmarking for various tolerances $\tau$. Let $\mathcal{P}$ denote the test problem set. A point $(\alpha,y)$ on a curve in a data profile indicates that a fraction of $y|\mathcal{P}|$ problems were solved, to tolerance $\tau$, given $(n+1)\times\alpha$ function evaluations. We say an algorithm solves a problem to tolerance $\tau$ if it finds an iterate $x_{k}$ satisfying
@@
Algorithm $b$ step size $s_{0}\dagger$ $\epsilon$ $\theta$ $\sigma_{0}$ $\sigma_{\min}^{\star}$ samp. radius
DFQRM - - - $10^{-5}$ 00$\ddagger$0 ‡ $1$ $0.01$
ZORO $1$ $\frac{0.005}{s_{0}\log n}$ $0.1n$ $5\times 10^{-4}$
adaZORO $0.5$ $\frac{0.005}{s_{0}\log n}$ $0.1n$ $5\times 10^{-4}$
ZORO-FA $1$ $0.1n$ $0.01$ $0.25$ $\frac{1}{s_{0}\log n}$
Nelder-Mead - - - - - - $0.001$
Table 2: Parameters used in the Moré-Garbow-Hillstrom benchmarking experiments of Section 4.2. $\dagger$: Here $s_{0}$ refers to the target sparsity in ZORO, and the initial target sparsity of adaZORO and ZORO-FA (which both dynamically adjust this parameter). $n$ refers to the problem size. $\ddagger$: In DFQRM, $\theta$ has a slightly different meaning. $\star$: In Nelder-Mead this is the minimum simplex size.
@@ 4.3 Discussion
On functions exhibiting gradient sparsity ZORO-FA outperforms existing benchmarks. Even though ZORO and adaZORO are given the exact gradient sparsity parameter, the fact that they cannot decrease their step-size or sampling radius means that their progress eventually plateaus. For the MGH test functions ZORO and adaZORO struggle, while Nelder-Mead, DFQRM, and ZORO-FA perform well. It was surprising, at least to us, that Nelder-Mead performed so well. We note however that this performance came at high computational expense, with Nelder-Mead typically taking 10–100 times longer to run.
@@ 4.4 Hidden Gradient Compressibility
In contrast to Section 4.1, the MGH functions of Section 4.2 are not explicitly designed to possess sparse or compressible gradients. Nonetheless we observe that ZORO-FA works well, and is frequently able to make progress with $s_{k}\ll n$. To investigate this further, we plot the magnitudes of the entries of $\nabla f(x)$, sorted from largest to smallest, for two MGH test functions (with $n=500$) at $20$ randomly selected $x$. Letting $|\nabla f(x)|_{(j)}$ denote the $j$-th largest-in-magnitude component of $\nabla f(x)$, we plot the mean of $|\nabla f(x)|_{(j)}$ over all $20$ values of $x$, as well as the minimum and maximum values of $|\nabla f(x)|_{(j)}$ observed.
We also plot a trajectory of ZORO-FA applied to these two functions, as well as the sparsity levels found at each iteration ($s_{k}$). The results are displayed in Figure 3. As is clear, some MGH test functions (e.g., rosex) exhibit extreme gradient compressibility, while others (e.g., trig) do not. As suggested above, ZORO-FA can indeed make progress using $s_{k}$ smaller than $n$. It is interesting that ZORO-FA uses lower values of $s_{k}$ for the function (trig) whose gradients appear less compressible. We suspect this might be due to the interplay of the sparsity and Lipschitz continuity of the gradients. As ZORO-FA searches for a suitable sparsity level ($s_{j}$) and Lipschitz constant ($\sigma_{j}$) simultaneously, it might ‘miss out’ on gradient sparsity when the Lipschitz constant is large. This further emphasizes the need for further research into optimization methods which can take advantage of the presence of gradient sparsity, yet are robust to its absence.
Figure 3: Top row: Gradient magnitude profiles for two MGH functions: rosex (Left), and trig (Right). The solid line represents the mean (over 20 trials), while the shading represents the minimum and maximum values of gradient component magnitudes observed. Bottom row: The successful sparsity level for ZORO-FA (left axes) as well as the objective function values (right axes) for rosex (Left), and trig (Right).